Golf-course designers have become concerned that old courses are becoming obsolete since new technology has given golfers the ability to hit the ball so far. Designers, therefore, have proposed that new golf courses need to be built expecting that the average golfer can hit the ball more than 250 yards on average. Suppose a random sample of 198 golfers be chosen so that their mean driving distance is 248.8 yards. The population standard deviation is 47.3. Use a 5% significance level.

Calculate the followings for a hypothesis test where H0:μ=250 and H1:μ<250:

(a)    The test statistic is    
(b)    The P-Value is

The final conclustion is

A. There is not sufficient evidence to warrant rejection of the claim that the mean driving distance is equal to 250
B. There is sufficient evidence to warrant rejection of the claim that the mean driving distance is equal to 250
Answer #1

(a) The test statistic is  

Z = (Xbar - µ)/[σ/sqrt(n)]

We are given

Xbar = 248.8

µ = 250

σ = 47.3

n = 198

Z = (248.8 - 250)/[47.3/sqrt(198)]

Z = (248.8 - 250)/ 3.3615

Z = -0.3570

The test statistic is -0.3570.

(b) The P-Value is 0.3606

(by using z-table)

Here, P-value > α = 0.05

So, we do not reject the null hypothesis

The final conclusion is

A. There is not sufficient evidence to warrant rejection of the claim that the mean driving distance is equal to 250.